OpenStudy (anonymous):

Find the Derivative of F when: F(x)= (f(x)+2x)/(f(x)-2x)

OpenStudy (anonymous):

since you are not given f, your answer must have an f' in it correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then you need the quotient rule for this. \[(\frac{h}{k})'=\frac{kh'-hk'}{k^2}\] with \[h(x)=f(x)+2x, h'(x)=f'(x)+2, k(x)=f(x)-2x, k'(x)=f'(x)-2\]

OpenStudy (anonymous):

so your denominator will be \[(f(x)-2x)^2\]

OpenStudy (anonymous):

numerator will be \[(f(x)-2)\times (f'(x)+2)-(f(x)+2)\times (f'(x)-2)\] then algebra to clean it up

OpenStudy (anonymous):

with quotient rule i get (f(x)-2x)(f'x+2)-(f(x)+2x))(f'x-2))/(f(x)-2))^2 but it want it simplified and im having trouble

OpenStudy (anonymous):

make your life easy and replace \[f(x)=y, f'(x)=y'\] should make algebra look simpler

OpenStudy (anonymous):

alright thanks again!

OpenStudy (anonymous):

for my answer choices its factoring out either a 2 or a 4 in the numerator

OpenStudy (anonymous):

\[(y-2x)(y'+2)-(y+2x)(y'-2)\] \[yy'+2y-2xy'-4x-(yy'-2y+2xy'-4x)\]

OpenStudy (anonymous):

\[yy'+2y-2xy'-4x-yy'+2y-2xy'+4x\]

OpenStudy (anonymous):

looks like we are left with nothing but \[4y-4xy'\]

OpenStudy (anonymous):

or if you prefer \[4(y-xy)'\] or even \[4(f(x)-xf'(x))\]

OpenStudy (anonymous):

ok i think i got it now

OpenStudy (anonymous):

check my algebra because i am doing it as i type, so i might have made an error, but i think it is right

OpenStudy (anonymous):

would you mind helping me with one last one?

OpenStudy (anonymous):

go ahead, i will help if i can

OpenStudy (anonymous):

find the derivative of: y=(cosx-1)/(sinx)

OpenStudy (anonymous):

this should be ok

OpenStudy (anonymous):

I derived it but cant simplify it in terms of their answer choices

OpenStudy (anonymous):

\[y'=\frac{\sin(x)\times (-\sin(x))-(\cos(x)-1)\times \cos(x))}{\sin^2(x)}\] is a start

OpenStudy (anonymous):

now we work in the numerator and get \[-\sin^2(x)-\cos^2(x)+\cos(x)\] if i am not mistaken

OpenStudy (anonymous):

and since \[\sin^2(x)+\cos^2(x)=1\] we have \[-\sin^2(x)-\cos^2(x)=-1\] so it looks like we get \[\frac{\cos(x)-1}{\sin^2(x)}\]

OpenStudy (anonymous):

we can go further i think

OpenStudy (anonymous):

I got up to that point but all of my answer choices have a 1 in the numerator

OpenStudy (anonymous):

since \[\sin^2(x)=1-\cos^2(x)=(1-\cos(x))(1+\cos(x))\]

OpenStudy (anonymous):

so we have \[\frac{\cos(x)-1}{(1-\cos(x))(1+\cos(x))}\] and \[\frac{\cos(x)-1}{1-\cos(x)}=-1\] we get \[\frac{-1}{1+\cos(x)}\] or maybe \[-\frac{1}{1+\cos(x)}\]

OpenStudy (anonymous):

Thank you so much, you were very helpful!

OpenStudy (anonymous):

there are many ways you can write this yw

OpenStudy (anonymous):

my big problem is sometimes I don't recognize some of the trig identities and it makes it very difficult to simplify. Thank You.